a*b=-a-b-2ab; a⊕b=3a+3b


1
Expert's answer
2021-12-13T14:05:03-0500

Let the operations * and ⊕ be defined on the set of integers.

The operation * on "\Z" is associative, if for every "a, b, c, ∈ \Z," we have

"(a ^* b) ^* c = a^* (b^*c)."

We have


"(a ^* b) ^* c=(-a-b-2ab) ^* c"

"=a+b+2ab-c+2(a+b+2ab)c"

"=a+b-c+2ab+2ac+2bc+4abc"

"a^* (b^*c)=a^* (-b-c-2bc)"

"=-a+b+c+2bc+2a(b+c+2bc)"

"=-a+b+c+2ab+2ac+2bc+4abc"

Suppose "(a ^* b) ^* c = a^* (b^*c), a, b, c \in \Z." Then


"a+b-c+2ab+2ac+2bc+4abc"

"=-a+b+c+2ab+2ac+2bc+4abc"

"=>a=c"

Hence the statement "(a ^* b) ^* c = a^* (b^*c)" is False for "a\not=c, a, b,c \in \Z."

The operation "^*" is not associative.


The operation "^*" on "\Z" is commutative, if for every "a, b, ∈ \Z," we have

"a ^* b = b^* a"

We have


"a ^* b =-a-b-2ab"

"b ^* a =-b-a-2ba=-a-b-2ab"

Hence the statement "a ^* b = b^* a" is True for "a, b \in \Z."

The operation "^*" is commutative.


The operation "^*" on "\Z" is left-distributive over "\u2295" , if for every "a, b, c∈ \Z," we have "a ^*( b\u2295c) =(a^*b)\u2295(a^*c)"

We have


"a ^*( b\u2295c)=-a-(3b+3c)-2a(3b+3c)"

"=-a-3b-3c-6ab-6ac"

"(a^*b)\u2295(a^*c) =3(-a-b-2ab)+3(-a-c-2ac)"


"=-6a-3b-3c-6ab-6ac"

Suppose "a ^*( b⊕c) =(a^*b)⊕(a^*c) , a, b, c \in \Z." Then


"-a-3b-3c-6ab-6ac"

"=-6a-3b-3c-6ab-6ac"

"=>a=0"

Hence the statement "a ^*( b\u2295c) =(a^*b)\u2295(a^*c)" is False for "a\not=0, a, b,c \in \Z."

The operation "^*" is not left-distributive over "\u2295."


The operation "^*" on "\Z" is right-distributive over "\u2295" , if for every "a, b, c∈ \Z," we have "( b\u2295c)^*a =(b^*a)\u2295(c^*a)"

We have


"( b\u2295c)^*a =-(3b+3c)-a-2a(3b+3c)"

"=-a-3b-3c-6ab-6ac"

"(b^*a)\u2295(c^*a) =3(-b-a-2ba)+3(-c-a-2ca)"

"=-6a-3b-3c-6ab-6ac"

Suppose "( b⊕c)^*a =(b^*a)⊕(c^*a) , a, b, c \in \Z." Then


"-a-3b-3c-6ab-6ac"

"=-6a-3b-3c-6ab-6ac"

"=>a=0"

Hence the statement "( b\u2295c)^*a =(b^*a)\u2295(c^*a)" is False for "a\not=0, a, b,c \in \Z."

The operation "^*" is not right-distributive over "\u2295."


The operation "\u2295" on "\Z" is associative, if for every "a, b, c, ∈ \Z," we have

"(a \u2295 b) \u2295 c = a\u2295(b\u2295c)."

We have


"(a\u2295 b) \u2295c=(3a+3b) \u2295c=3(3a+3b)+3c"

"=9a+9b+3c"

"a\u2295(b\u2295c)=a\u2295(3b+3c)=3a+3(3b+3c)"

"=3a+9b+9c"

Suppose "(a ⊕ b) ⊕ c = a⊕(b⊕c) , a, b, c \in \Z." Then

"9a+9b+3c=3a+9b+9c=>a=c"

Hence the statement "(a \u2295 b) \u2295 c = a\u2295(b\u2295c)" is False for "a\not=c, a, b,c \in \Z."

The operation "\u2295" is not associative


The operation "\u2295" on "\Z" is commutative, if for every "a, b, ∈ \Z," we have

"a\u2295 b = b\u2295 a"

We have


"a \u2295b =3a+3b"

"b\u2295a =3b+3a=3a+3b"

Then "a ⊕ b =3a+3b=3b+3a= b⊕a, a,b\in\Z."

The operation "\u2295" is commutative.


The operation "\u2295" on "\Z" is left-distributive over "^*" , if for every "a, b, c∈ \Z," we have "a \u2295( b^*c) =(a\u2295b)^*(a\u2295c)."

We have


"a \u2295( b^*c)=3a+3(-b-c-2bc)"

"=3a-3b-3c-6bc"

"(a\u2295b)^*(a\u2295c)=-(3a+3b)-(3a+3c)"


"-2(3a+3b)(3a+3c)"

"=-6a-3b-3c-18a^2-18ab-18ac-18bc"

Let "a=1, b=c=0." Then


"a \u2295( b^*c)=1 \u2295( 0^*0)"

"=3(1)-3(0)-3(0)-6(0)(0)=3"

"(a\u2295b)^*(a\u2295c)=(1\u22950)^*(1\u22950)"

"=-6(1)-3(0)-3(0)-18(1)^2"

"-18(1)(0)-18(1)(0)-18(0)(0)=-24"

Since "3\not=-24," the operation "\u2295" is not left-distributive over "^*."


The operation "\u2295" on "\Z" is right-distributive over "^*" , if for every "a, b, c∈ \Z," we have "( b^*c) \u2295a=(b\u2295a)^*(c\u2295a)."

We have


"( b^*c) \u2295a=3(-b-c-2bc)+3a"

"=3a-3b-3c-6bc"

"(b\u2295a)^*(c\u2295a)=-(3b+3a)-(3c+3a)"

"-2(3b+3a)(3c+3a)"

"=-6a-3b-3c-18bc-18ac-18ab-18a^2"

Let "a=1, b=c=0." Then


"( b^*c) \u2295a=( 0^*0) \u22951=3(1)-3(0)-3(0)"

"-6(0)(0)=3"

"(b\u2295a)^*(c\u2295a)=(0\u22951)^*(0\u22951)"

"=-6(1)-3(0)-3(0)-18(0)(0)-18(1)(0)"

"-18(1)(0)-18(1)^2=-24"

Since "3\not=-24," the operation "\u2295" is not right-distributive over "^*."